Evaluate \oint_{C} \left ( -2xy+y^{2} \right )dx+x^{2}dy where C is the boundary of the graphs of y=2x^{2} from (0,0) to (2,8) followed by a straight line segment from (2,8) to (0,0).
Answer #1

y (2,8) 8 y=42 2 Y22x2 =SS (4х-24) dy de- = 1 (424 8 dn = The boundary c is shown below, § (-2xy + y2) dx + x dy p=-209+ y; Q